OpenStudy (anonymous):

A card is drawn from a standard deck. What is the probability that the card is and ace (A) or a Heart (H)?

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

P (A) =4/52 =1/13 P (H) = 13/52 =1/4 P (A or H) [additive] = 17/52 P (A or H) [exclusive] = 16/52 = 4/13

OpenStudy (unklerhaukus):

the difference between additive and exclusive is the Ace of hearts, is this a ace or a heart ( you could say yes, or no depending on what you mean by 'or')